# Multimodal 3D Genome Pre-training

### v1

@@ Abstract
Deep learning techniques have driven significant progress in various analytical tasks within 3D genomics in computational biology. However, a holistic understanding of 3D genomics knowledge remains underexplored. Here, we propose MIX-HIC, the first multimodal foundation model of 3D genome that integrates both 3D genome structure and epigenomic tracks, which obtains unified and comprehensive semantics. For accurate heterogeneous semantic fusion, we design the cross-modal interaction and mapping blocks for robust unified representation, yielding the accurate aggregation of 3D genome knowledge. Besides, we introduce the first large-scale dataset comprising over 1 million pairwise samples of Hi-C contact maps and epigenomic tracks for high-quality pre-training, enabling the exploration of functional implications in 3D genomics. Extensive experiments show that MIX-HIC can significantly surpass existing state-of-the-art methods in diverse downstream tasks. This work provides a valuable resource for advancing 3D genomics research.
Machine Learning, ICML
@@ 1 Introduction
The three-dimensional (3D) organization of chromosomes within the nucleus plays a pivotal role in gene regulation and cellular function (Zhang et al., 2024; Monteagudo-Sánchez et al., 2024). Key topological features of 3D genome, such as chromatin loops, are essential for cell-type-specific transcriptional regulation. High-resolution 3D chromatin interactions can be quantified through high-throughput chromosome conformation capture (Hi-C) technique (Lieberman-Aiden et al., 2009). Understanding the mechanisms of how the 3D genome influences gene expression can unveil pivotal insights into cellular functionality, developmental biology, and disease mechanisms (Ferrer & Dimitrova, 2024).
@@
Recently, computational models have emerged as a powerful tool to unravel the intricate associations between the 3D chromatin structure, epigenome, and transcriptome. Existing approaches predict various genomic features, including 3D chromatin contact maps (Yang et al., 2023; Tan et al., 2023; Fudenberg et al., 2020), chromatin loops (Salameh et al., 2020; Matthey-Doret et al., 2020; Wang et al., 2022), and gene expression (Zhang et al., 2023; Karbalayghareh et al., 2022), often leveraging DNA sequences, Hi-C contact maps and epigenomic tracks. Though success, most of these methods are limited to a single specific task and struggle to integrate the diverse and heterogeneous information of the 3D genome, hindering a comprehensive understanding of its complex organization.
@@
Recent progress in large-scale foundation models have demonstrated remarkable success in various fields of computations biology, such as molecular representations (Feng et al., 2024; Chang & Ye, 2024), proteomics (Zhao et al., 2024; Gong et al., 2024), and genomics (Li et al., 2024; Schiff et al., 2024). Inspired by these advancements, we aim to develop a multimodal foundation model to address the above-mentioned limitations of analyzing 3D genomic downstream tasks in isolation.
@@
Modeling the multimodal foundation model of 3D genome introduces three key challenges. First, Hi-C contact maps and epigenomic tracks have inherently distinct characteristics, making integration difficult. Simply aligning features from the two modalities and projecting them into a unified latent space would primarily capture modal-invariant knowledge like gene regulatory mechanisms, which are governed by both chromatin spatial organization and epigenomic tracks. However, this approach tends to overlook modal-specific characteristics, such as precise chemical modifications and chromatin states revealed by epigenomic tracks, which are essential factors for fine-grained 3D genome analysis. This can lead to information loss and degrade downstream task performance (Refer to Appendix B for theoretical analysis). Second, the unified representation from heterogeneous 3D genomic and epigenomic data must exhibit robust generalization capabilities, otherwise, the model may struggle to adapt effectively to diverse downstream tasks e.g., generation and regression. Third, uncovering implicit semantic relationships between 3D genomic and epigenomic data is crucial for addressing the data scarcity problem in 3D genomics. Especially, the high experimental costs of Hi-C sequencing in real-world applications limit data accessibility, leading to incomplete representations and degraded model performance. Pre-training a model to learn implicit multimodal structural connections offers significant benefits for downstream tasks, particularly when only single-modality data is available. This enables the model to leverage structural knowledge to compensate for missing modality semantics, thereby enriching the overall data representation.
@@
Hence, we introduce MIX-HIC, the first multimodal foundation model for 3D genomics to extract fine-grained knowledge from 3D genome and epigenomic tracks, enabling efficient adaptation to diverse downstream tasks with superior performance. Specifically, MIX-HIC first incorporates two distinct encoders to capture the refined features from 3D genome contact maps and epigenomic tracks, leveraging cell type-specific information for accurate predictions in novel cell types. To address the challenge of integrating heterogeneous data, we propose a cross-modal interaction block to capture both modal-invariant and modal-specific representations, regularized by contrastive learning and orthogonal constraints, preserving both shared and distinctive information across modalities. Additionally, a cross-modal mapping block facilitates information exchange between modalities, ensuring robust representation even with single-modality input. To comprehensively capture 3D genome knowledge, we have curated a large-scale dataset which consists of 1,275,948 pair samples of Hi-C contact maps and epigenomic tracks for rapid adaptation to downstream tasks through task-specific decoders. Notably, this is the largest paired dataset for the 3D genome analysis to date. MIX-HIC is evaluated across diverse downstream tasks, demonstrating its effectiveness and robustness in comparison to other state-of-the-art methods. In summary, the main contributions are:
@@
  • For holistic representation learning for 3D genome analysis, we present the largest paired dataset of Hi-C and epigenomic tracks, comprising over 1 million samples.
@@ 2 Related Works
This work evaluates the effectiveness of MIX-HIC on three downstream tasks, including Hi-C contact maps prediction, chromatin loops prediction, and CAGE-seq expression prediction.
Existing methods for Hi-C contact maps prediction from DNA sequences show promise but lack cross-cell-type generalization (Fudenberg et al., 2020; Schwessinger et al., 2020). Models like EPCOT (Zhang et al., 2023) and C.Origami (Tan et al., 2023) improve this by integrating DNA sequence with cell-type-specific epigenomic tracks. Epiphany (Yang et al., 2023) offers a more efficient solution using only the epigenomic tracks.
Chromatin loop prediction methods are broadly divided into statistical and supervised learning methods. Statistical methods like HiCExplorer (Wolff et al., 2020) and ChromoSight (Matthey-Doret et al., 2020) rely on contact frequency distributions or expert-defined templates to identify loops. Supervised learning methods, including Peakachu (Salameh et al., 2020), DLoopCaller (Wang et al., 2022) leverage labeled data and sophisticated architectures for loop prediction. RefHiC (Zhang & Blanchette, 2023) employs a coarse-to-fine training strategy that integrates multi-resolution Hi-C contact maps through contrastive learning mechanisms for model pre-training. However, RefHiC is limited by its reliance on small-scale data semantics, which hinders its generalizability to other downstream tasks in 3D genome.
CAGE-seq expression prediction typically uses multimodal inputs, including DNA sequences, epigenomic tracks, and Hi-C contact maps. Enformer (Avsec et al., 2021) employs transformers for modeling DNA sequences, while EPCOT (Zhang et al., 2023) integrates DNA sequences and epigenomic tracks with transformers or long short-term memory (LSTM) (Hochreiter, 1997). GraphReg (Karbalayghareh et al., 2022) integrates epigenomic tracks and Hi-C contact maps using graph attention networks for expression prediction.
@@ 2.2 Foundation Models in Computational Biology
The emergence of foundation models has significantly advanced various fields of computational biology. For example, EvoRank (Zhao et al., 2024) has demonstrated remarkable capabilities by harnessing extensive protein sequence datasets to derive latent representations through sequence alignment. VQDNA (Li et al., 2024) employs large-scale DNA sequences to learn adaptive tokenization through vector-quantized codebooks. To capture more comprehensive views of data, multimodal foundation models such as ESM-IF (Hsu et al., 2022) combine protein sequences and 3D structures to learn functional and structural representations, advancing protein function design and prediction. Similarly, UniCorn (Feng et al., 2024) integrates 2D and 3D molecular views through contrastive learning, effectively capturing complementary bimodal features. These foundation models have showcased strong representation learning capabilities of foundation models, inspiring the development of a universal model capable of effectively addressing various downstream tasks in the 3D genomic field.
@@ 3 Data Processing and Preparation
To distill the comprehensive semantics from the 3D genome, we collect and refine a large-scale dataset for pre-training MIX-HIC, using publicly available data from the hg38 assembly. The Hi-C contact maps are obtained from the 4DN Data Portal [^1] https://data.4dnucleome.org/, while the epigenomic tracks (ATAC-seq and DNase-seq), CAGE-seq expression data, and CTCF ChIA-PET (Li et al., 2010) chromatin loops are downloaded from the ENCODE Portal  [^2] https://www.encodeproject.org/. MIX-HIC processes 250,000 base pair (bp) genomic windows as input. The Hi-C contact maps are represented as $50\times 50$ matrices per window at 5 kb resolution, while epigenomic tracks are averaged over every 100 bps to reduce data variability, generating 2,500-length sequences. Each bin in the Hi-C matrix corresponds to specific $x$ and $y$ coordinates, representing two distinct genomic segments. Their respective epigenomic sequences are concatenated into a 5,000-length representation for that bin. Thus, Hi-C contact maps and epigenomic tracks are treated as images and sequences, respectively. We focus on four cell lines for pre-training, including HepG2, HCT116, IMR90, and WTC111. To ensure data quality, we filter out windows with fewer than 10% non-zero Hi-C interactions or insufficient contact signals. This process results in a total of 1,275,948 pairwise samples of Hi-C contact maps and epigenomic tracks within chromatin loop regions for pre-training. Table 1 summarizes the sample counts before and after this filtering process. Further details on data processing, reference number, and downstream task data are provided in Appendix A.1.
@@ 4.1 Self-supervised Pre-training
Self-supervised learning (Chen et al., 2020) offers a powerful solution for learning unified and comprehensive representations from heterogeneous 3D genomic and epigenomic data. By leveraging large-scale pairwise data during pre-training, MIX-HIC effectively captures inherent biological patterns and relationships across these bimodal data, significantly enhancing its adaptability to diverse downstream tasks. As shown in Figure 1, the pre-training phase includes a feature extraction block, cross-modal interaction block, and cross-modal mapping block. The feature extraction block employs specialized encoders for epigenomic tracks and Hi-C contact maps, respectively, to generate refined representations. The learned representations are then fed into the cross-modal interaction block to learn both modal-specific and modal-invariant features. This learning process is regularized through a combination of contrastive learning loss and orthogonal loss. Finally, a cross-modal mapping block further explores latent connections and complementary information between the two modalities.
@@
Feature extraction block. The feature extraction block consists of Hi-C and epigenomic feature encoders. Given the Transformer’s capability to model long-range dependencies and its flexibility in capturing both spatial interactions (Hi-C contact maps) and sequential relationships (epigenomic tracks), we utilize a Transformer-based architecture for robust feature extraction and multimodal integration. Similar to Vision Transformer (ViT) (Dosovitskiy et al., 2021), the single-channel Hi-C contact map $X_{M}\in\mathbb{R}^{1\times H\times W}$ is first transformed into a sequence of flattened patches $X_{M}^{p}\in\mathbb{R}^{N\times D}$. Here, $H$ and $W$ are the height and width of the Hi-C contact map, both equal to 50. The dimension $D=1\times P\times P$ corresponds to the initial size of each patch, where the patch size $P$ is set to 2. The total number of these patches $N=H\times W/P^{2}$ is the resulting number of patches, which becomes the input sequence length. In the Hi-C feature encoder, a feedforward network projects these patches into an embedding $X_{M}^{0}\in\mathbb{R}^{N\times C}$, where $C$ denotes the predefined feature dimension. This embedding is then refined through three cascaded encoder layers, where each layer consists of $T$ Transformer blocks followed by a downsampling layer, yielding progressively refined embeddings $X_{M}^{i}\in\mathbb{R}^{\alpha_{i}\times C_{i}}$, where the sequence length $\alpha_{i}$ is defined as $\alpha_{i}=\frac{N}{4^{i}}$ and the feature dimension $C_{i}$ grows exponentially as $C_{i}=2^{i}C$, with the encoder layer $i\in\{1,2,3\}$. Finally, a bottleneck layer equipped with $T$ Transformer blocks is applied to produce the Hi-C contact map embedding $X_{M}^{B}\in\mathbb{R}^{\alpha_{3}\times C_{3}}$.
@@
The epigenomic encoder processes input sequences $X_{E}\in\mathbb{R}^{L_{1}\times O}$, where $L_{1}=5,000$ is the initial sequence length, and $O$ represents the two epigenomic tracks: ATAC-seq and DNase-seq. Convolutional layers with max-pooling operations extract an initial embedding $X_{E}^{0}\in\mathbb{R}^{L_{2}\times C}$, where $L_{2}=100$ corresponds to two genomic segments along both $x$ and $y$ axes in the Hi-C contact map. Similarly, the embedding $X_{E}^{0}$ is processed through three encoder layers, where each layer consists of $T$ Transformer blocks followed by a downsampling layer, resulting in $X_{E}^{i}\in\mathbb{R}^{\beta_{i}\times C_{i}}$, where the sequence length $\beta_{i}=\frac{L_{2}}{2^{i}}$ for $i\in\{1,2,3\}$. Then, the refined epigenomic representation $X_{E}^{\text{B}}\in\mathbb{R}^{\beta_{3}\times C_{3}}$ is also derived from a final bottleneck layer.
@@
Cross-modal interaction block. Conventional multimodal learning architectures (Yang et al., 2024) often utilize contrastive learning to project features from different modalities into a shared embedding space. However, these modalities inherently contain both homogeneous and heterogeneous information. Direct aligning all features risks losing essential modal-specific characteristics, potentially diminishing downstream task performance (Dong et al., 2023; Wang et al., 2023). This is further analyzed in the Appendix B.
@@
To address this, our cross-modal interaction block captures both modal-specific and modal-invariant representations, enabling a more comprehensive understanding of the data. Specifically, we employ four independent dense networks to process the Hi-C contact map representation $X_{M}^{B}$ and the epigenomic representation $X_{E}^{B}$. This generates the condensed modal-invariant representations $X_{M}^{I}\in\mathbb{R}^{\alpha_{3}\times C_{2}}$ and $X_{E}^{I}\in\mathbb{R}^{\beta_{3}\times C_{2}}$, as well as the modal-specific representations $X_{M}^{S}\in\mathbb{R}^{\alpha_{3}\times C_{2}}$ and $X_{E}^{S}\in\mathbb{R}^{\beta_{3}\times C_{2}}$. The mean pooling operation is then applied along the sequence length dimension, producing $\hat{X_{M}^{I}}$, $\hat{X_{E}^{I}}$, $\hat{X_{M}^{S}}$, and $\hat{X_{E}^{S}}$, each with a feature dimension of $C_{2}$. To ensure that the modal-invariant features capture shared knowledge across modalities, we incorporate a contrastive learning loss to regularize these representations. We propose a unified contrastive loss function $\mathcal{L}_{\text{pair}}$ to compute the similarity between two modalities $\mathcal{A}$ and $\mathcal{B}$ as follows:

### v2

@@ Abstract
Deep learning techniques have driven significant progress in various analytical tasks within 3D genomics in computational biology. However, a holistic understanding of 3D genomics knowledge remains underexplored. Here, we propose MIX-HIC, the first multimodal foundation model of 3D genome that integrates both Hi-C contact maps and epigenomic tracks, which obtains unified and comprehensive semantics. For accurate heterogeneous semantic fusion, we design the cross-modal interaction and mapping blocks for robust unified representation, yielding the accurate aggregation of 3D genome knowledge. Besides, we introduce the first large-scale dataset comprising over 1 million pairwise samples of Hi-C contact maps and epigenomic tracks for high-quality pre-training, enabling the exploration of functional implications in 3D genomics. Extensive experiments show that MIX-HIC significantly surpasses existing state-of-the-art methods in diverse downstream tasks. This work provides a valuable resource for advancing 3D genomics research.
**footnotetext: Correspondence to Zhi-An Huang <huang.za@cityu-dg.edu.cn>, Hui Xiong <xionghui@ust.hk>.footnotetext: The source code of MIX-HIC is available at https://github.com/myang998/MIX-HIC.
@@ 1 Introduction
The three-dimensional (3D) organization of chromosomes within the nucleus plays a pivotal role in gene regulation and cellular function zhang2024computational ; monteagudo2024impact . Key topological features of the 3D genome, such as chromatin loops that bring distant regulatory elements into close physical proximity with their target genes, are essential for cell-type-specific transcriptional regulation. High-resolution 3D chromatin interactions can be quantified through high-throughput chromosome conformation capture (Hi-C) technique lieberman2009comprehensive . Understanding the mechanisms of how the 3D genome influences gene expression can unveil pivotal insights into cellular functionality, developmental biology, and disease mechanisms ferrer2024transcription .
@@
Recently, computational models have emerged as a powerful tool to unravel the intricate associations between the 3D chromatin structure, epigenome, and transcriptome. Existing approaches predict various genomic features, including 3D chromatin contact maps yang2023epiphany ; tan2023cell ; fudenberg2020predicting , chromatin loops salameh2020supervised ; matthey2020computer ; wang2022dloopcaller , and gene expression zhang2023generalizable ; karbalayghareh2022chromatin , often leveraging DNA sequences, Hi-C contact maps and epigenomic tracks. Though successful, most of these methods are limited to a single specific task and struggle to integrate the diverse and heterogeneous information of the 3D genome, hindering a comprehensive understanding of its complex organization.
@@
Recent progress in large-scale foundation models has demonstrated remarkable success in various fields of computational biology, such as molecular representations wang2024exploring ; chang2024bidirectional , medical imaging li2023llava ; xia2024cares , proteomics cheng2024training ; li2024prosst , and genomics icml2024vqdna ; schiff2024caduceus . Inspired by these advancements, we aim to develop a multimodal foundation model to address the above-mentioned limitations of analyzing 3D genomic downstream tasks in isolation.
@@
Modeling the multimodal foundation model of 3D genome introduces three key challenges. First, Hi-C contact maps and epigenomic tracks have inherently distinct characteristics, making integration difficult. Simply aligning features from the two modalities and projecting them into a unified latent space would primarily capture modal-invariant knowledge like gene regulatory mechanisms, which are governed by both chromatin spatial organization and epigenomic tracks. However, this approach tends to overlook modal-specific characteristics, such as precise chemical modifications and chromatin states revealed by epigenomic tracks, which are essential factors for fine-grained 3D genome analysis. This can lead to information loss and degrade downstream task performance (Refer to Appendix B for theoretical analysis). Second, the unified representation from heterogeneous 3D genomic and epigenomic data must exhibit robust generalization capabilities; otherwise, the model may struggle to adapt effectively to diverse downstream tasks e.g., generation and regression. Third, uncovering implicit semantic relationships between 3D genomic and epigenomic data is crucial for addressing the data scarcity problem in 3D genomics. Especially, the high experimental costs of Hi-C sequencing in real-world applications limit data accessibility, leading to incomplete representations and degraded model performance. Pre-training a model to learn implicit multimodal structural connections offers significant benefits for downstream tasks, particularly when only single-modality data is available. This enables the model to leverage structural knowledge to compensate for missing modality semantics, thereby enriching the overall data representation.
@@
Hence, we introduce MIX-HIC, the first multimodal foundation model for 3D genomics to extract fine-grained knowledge from 3D genome and epigenomic tracks, enabling efficient adaptation to diverse downstream tasks with superior performance. Specifically, MIX-HIC first incorporates two distinct encoders to capture the refined features from 3D genome contact maps and epigenomic tracks, leveraging cell type-specific information for accurate predictions in novel cell types. To address the challenge of integrating heterogeneous data, we propose a cross-modal interaction block to capture both modal-invariant and modal-specific representations, regularized by contrastive learning and orthogonal constraints, preserving both shared and distinctive information across modalities. Additionally, a cross-modal mapping block facilitates information exchange between modalities, ensuring robust representation even with single-modality input. To comprehensively capture 3D genome knowledge, we have curated a large-scale dataset that consists of 1,275,948 pair samples of Hi-C contact maps and epigenomic tracks for rapid adaptation to downstream tasks through task-specific decoders. Notably, this is the largest paired dataset for the 3D genome analysis to date. MIX-HIC is evaluated across diverse downstream tasks, demonstrating its effectiveness and robustness in comparison to other state-of-the-art methods. In summary, the main contributions are:
@@
  • For holistic representation learning in 3D genome analysis, we present the largest paired dataset of Hi-C and epigenomic tracks, comprising over 1 million samples.
@@ 2 Related Works
This work evaluates the effectiveness of MIX-HIC on three downstream tasks, including Hi-C contact map prediction, chromatin loop detection, and CAGE-seq expression prediction.
Existing methods for Hi-C contact map prediction from DNA sequences show promise but lack cross-cell-type generalization fudenberg2020predicting ; schwessinger2020deepc . Models like EPCOT zhang2023generalizable and C.Origami tan2023cell improve this by integrating DNA sequence with cell-type-specific epigenomic tracks. Epiphany yang2023epiphany offers a more efficient solution using only the epigenomic tracks.
Chromatin loop detection methods are broadly divided into statistical and supervised learning methods. Statistical methods like HiCExplorer wolff2020galaxy and ChromoSight matthey2020computer rely on contact frequency distributions or expert-defined templates to identify loops. Supervised learning methods, including Peakachu salameh2020supervised , DLoopCaller wang2022dloopcaller leverage labeled data and sophisticated architectures for loop detection. RefHiC zhang2023reference employs a coarse-to-fine training strategy that integrates multi-resolution Hi-C contact maps through contrastive learning mechanisms for model pre-training. However, RefHiC is limited by its reliance on small-scale data semantics, which hinders its generalizability to other downstream tasks.
CAGE-seq expression prediction typically uses multimodal inputs, including DNA sequences, epigenomic tracks, and Hi-C contact maps. Enformer avsec2021effective employs transformers for modeling DNA sequences, while EPCOT zhang2023generalizable integrates DNA sequences and epigenomic tracks with transformers or long short-term memory (LSTM) hochreiter1997long . GraphReg karbalayghareh2022chromatin integrates epigenomic tracks and Hi-C contact maps using graph attention networks for expression prediction.
@@ 2.2 Foundation Models in Computational Biology
The emergence of foundation models has significantly advanced various fields of computational biology. For example, EvoRank zhuang2024pre has demonstrated remarkable capabilities by harnessing extensive protein sequence datasets to derive latent representations through sequence alignment. VQDNA icml2024vqdna employs large-scale DNA sequences to learn adaptive tokenization through vector-quantized codebooks. To capture more comprehensive views of data, multimodal foundation models such as ESM-IF hsu2022learning combine protein sequences and 3D structures to learn functional and structural representations, advancing protein function design and prediction. Similarly, UniCorn feng2024unicorn integrates 2D and 3D molecular views through contrastive learning, effectively capturing complementary bimodal features. These foundation models have showcased strong representation learning capabilities of foundation models, inspiring the development of a universal model capable of effectively addressing various downstream tasks in the 3D genomic field.
@@ 3 Data Processing and Preparation
To distill the comprehensive semantics from the 3D genome, we collect and refine a large-scale dataset for pre-training MIX-HIC, using publicly available data from the hg38 assembly. The Hi-C contact maps are obtained from the 4DN Data Portal [^1] https://data.4dnucleome.org/, while the epigenomic tracks (ATAC-seq and DNase-seq, which measure how ‘open’ or accessible DNA is for transcription), CAGE-seq expression data (which directly quantifies gene activity levels), and CTCF ChIA-PET li2010chia chromatin loops (which identify high-confidence interactions mediated by the key architectural protein CTCF) are downloaded from the ENCODE Portal  [^2] https://www.encodeproject.org/. Due to the high cost of deep sequencing for Hi-C experiments chang2024droplet , publicly available datasets are typically limited to resolutions of 5 kb, 10 kb, or coarser. A 5 kb resolution is a fine-grained and effective choice for deep learning models zhang2023generalizable ; karbalayghareh2022chromatin . The MIX-HIC model processes 250,000 base pair (bp) genomic windows, a size selected to encompass key regulatory structures like chromatin loops, ensuring most functional units are fully contained within the inputs yang2023epiphany ; salameh2020supervised . This results in Hi-C contact maps being represented as $50\times 50$ matrices at the 5 kb resolution. To reduce data variability, epigenomic tracks are averaged over every 100 bps, which generates 2,500-length sequences. Each bin in the Hi-C matrix corresponds to specific $x$ and $y$ coordinates, representing two distinct genomic segments. Their respective epigenomic sequences are concatenated into a 5,000-length representation for that bin. Thus, Hi-C contact maps and epigenomic tracks are treated as images and sequences, respectively. We focus on four cell lines for pre-training, including HepG2, HCT116, IMR90, and WTC111. To ensure data quality, we filter out windows with fewer than 10% non-zero Hi-C interactions or insufficient contact signals, which is a standard quality control step to remove uninformative data salameh2020supervised ; wang2022dloopcaller . Hi-C contact maps are inherently sparse, especially for long-range interactions distant from the diagonal. Including these extremely sparse, low-signal windows would introduce noise and degrade the model’s training process. As summarized in Table 1, this filtering process removed approximately 30% of the raw windows. Crucially, a massive and high-quality dataset of over 1.2 million sample pairs was retained, which is more than sufficient for robust pre-training. Further details on data processing, reference number, and downstream task data are provided in Appendix A.1.
@@ 4.1 Self-supervised Pre-training
Self-supervised learning simclr ; wang2020enaet ; wang2022importance offers a powerful solution for learning unified and comprehensive representations from heterogeneous 3D genomic and epigenomic data. By leveraging large-scale pairwise data during pre-training, MIX-HIC effectively captures inherent biological patterns and relationships across these bimodal data, significantly enhancing its adaptability to diverse downstream tasks. As shown in Figure 1, the pre-training phase includes a feature extraction block, cross-modal interaction block, and cross-modal mapping block. The feature extraction block employs specialized encoders for epigenomic tracks and Hi-C contact maps, respectively, to generate refined representations. The learned representations are then fed into the cross-modal interaction block to learn both modal-specific and modal-invariant features. This learning process is regularized through a combination of contrastive learning loss and orthogonal loss. Finally, a cross-modal mapping block further explores latent connections and complementary information between the two modalities.
@@
Feature extraction block. The feature extraction block consists of Hi-C and epigenomic feature encoders. Given the Transformer’s capability to model long-range dependencies and its flexibility in capturing both spatial interactions (Hi-C contact maps) and sequential relationships (epigenomic tracks), we utilize a Transformer-based architecture for robust feature extraction and multimodal integration. Similar to Vision Transformer (ViT) dosovitskiy2020image , the single-channel Hi-C contact map $X_{M}\in\mathbb{R}^{1\times H\times W}$ is first transformed into a sequence of flattened patches $X_{M}^{p}\in\mathbb{R}^{N\times D}$. Here, $H$ and $W$ are the height and width of the Hi-C contact map, both equal to 50. The dimension $D=1\times P\times P$ corresponds to the initial size of each patch, where the patch size $P$ is set to 2. The total number of these patches $N=H\times W/P^{2}$ is the resulting number of patches, which becomes the input sequence length. In the Hi-C feature encoder, a feedforward network projects these patches into an embedding $X_{M}^{0}\in\mathbb{R}^{N\times C}$, where $C$ denotes the predefined feature dimension. This embedding is then refined through three cascaded encoder layers, where each layer consists of $T$ Transformer blocks followed by a downsampling layer, yielding progressively refined embeddings $X_{M}^{i}\in\mathbb{R}^{\alpha_{i}\times C_{i}}$, where the sequence length $\alpha_{i}$ is defined as $\alpha_{i}=\frac{N}{4^{i}}$ and the feature dimension $C_{i}$ grows exponentially as $C_{i}=2^{i}C$, with the encoder layer $i\in\{1,2,3\}$. Finally, a bottleneck layer equipped with $T$ Transformer blocks is applied to produce the Hi-C contact map embedding $X_{M}^{B}\in\mathbb{R}^{\alpha_{3}\times C_{3}}$.
@@
The epigenomic encoder processes input sequences $X_{E}\in\mathbb{R}^{L_{1}\times O}$, where $L_{1}=5,000$ is the initial sequence length, and $O$ represents the two epigenomic tracks: ATAC-seq and DNase-seq. Convolutional layers with max-pooling operations extract an initial embedding $X_{E}^{0}\in\mathbb{R}^{L_{2}\times C}$, where $L_{2}=100$ corresponds to two genomic segments along both $x$ and $y$ axes in the Hi-C contact map. Note that processing epigenomic tracks at high resolution (e.g., 100bp) is common practice to avoid over-smoothing yang2023epiphany ; zhang2023generalizable . Similarly, the embedding $X_{E}^{0}$ is processed through three encoder layers, where each layer consists of $T$ Transformer blocks followed by a downsampling layer, resulting in $X_{E}^{i}\in\mathbb{R}^{\beta_{i}\times C_{i}}$, where the sequence length $\beta_{i}=\frac{L_{2}}{2^{i}}$ for $i\in\{1,2,3\}$. Then, the refined epigenomic representation $X_{E}^{\text{B}}\in\mathbb{R}^{\beta_{3}\times C_{3}}$ is also derived from a final bottleneck layer.
@@
Cross-modal interaction block. Conventional multimodal learning architectures yang2024employing often utilize contrastive learning to project features from different modalities into a shared embedding space. However, these modalities inherently contain both homogeneous and heterogeneous information. Direct aligning all features risks losing essential modal-specific characteristics, potentially diminishing downstream task performance dong2023simmmdg ; wang2023multi . This is further analyzed in Theorem 1 as follows.
Theorem 1. Let $\phi_{1}$ and $\phi_{2}$ be feature encoders for two modalities $z_{1}$ and $z_{2}$, respectively. If the encoded features $\mathbf{F}^{1}=\phi_{1}(z_{1})$ and $\mathbf{F}^{2}=\phi_{2}(z_{2})$ are perfectly aligned such that $\mathbf{F}^{1}=\mathbf{F}^{2}$, we have:
Remarks. The information gap $\Gamma_{q}:=\max\{U(z_{1};t),U(z_{2};t)\}-\min\{U(z_{1};t),U(z_{2};t)\}$ quantifies the effectiveness of the modalities in predicting target variable $t$, where $U(z_{j};t)$ represents the mutual information. Here, $\mathbb{E}$ represents the expectation, $q$ denotes the joint distribution of $(z_{1},z_{2},t)$, $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, and $h$ and $h^{\prime}$ are prediction functions for features and raw data, respectively. Theorem 1 demonstrates that perfect alignment results in prediction errors that are suboptimal by at least $\Gamma_{q}$ compared to using raw modalities directly. This information gap widens when information content is imbalanced across modalities. The complete proof is provided in Appendix B.
@@
To address this, our cross-modal interaction block captures both modal-specific and modal-invariant representations, enabling a more comprehensive understanding of the data. Specifically, we employ four independent dense networks to process the Hi-C contact map representation $X_{M}^{B}$ and the epigenomic representation $X_{E}^{B}$. This generates the condensed modal-invariant representations $X_{M}^{I}\in\mathbb{R}^{\alpha_{3}\times C_{2}}$ and $X_{E}^{I}\in\mathbb{R}^{\beta_{3}\times C_{2}}$, as well as the modal-specific representations $X_{M}^{S}\in\mathbb{R}^{\alpha_{3}\times C_{2}}$ and $X_{E}^{S}\in\mathbb{R}^{\beta_{3}\times C_{2}}$. The mean pooling operation is then applied along the sequence length dimension, producing $\hat{X_{M}^{I}}$, $\hat{X_{E}^{I}}$, $\hat{X_{M}^{S}}$, and $\hat{X_{E}^{S}}$, each with a feature dimension of $C_{2}$. To ensure that the modal-invariant features capture shared knowledge across modalities, we incorporate a contrastive learning loss to regularize these representations. We propose a unified contrastive loss function $\mathcal{L}_{\text{pair}}$ simclr to compute the similarity between two modalities $\mathcal{A}$ and $\mathcal{B}$ as follows: